A thin copper tube, 14 mm internal diameter and 0,5 mm wall thickness has closed ends and rotates at 1200 r/min while it is subjected to an internal gauge pressure of 222 Pa. Calculate the total hoop stress in the wall of the tube if the density of copper is 8900 kg/m3.


1
Expert's answer
2021-07-19T03:15:22-0400

"P_{abs} = P_g + P_{atm}\\
P_{abs} =222 + 8900*9.8*1\\
P_{abs} =87442N/m^2\\
\sigma_{h}= \frac{Pd}{2t}\\
\sigma_{h}= \frac{87442*(\frac{14+0.5}{1000})}{2*\frac{0.5}{1000}}\\
\sigma_{h}= 1267909 N/m^2"